# Center Biophysics and Biochemistry, (IVIC), Caracas, Venezuela

### v2

@@ Abstract
is obtained under randomization, this is does not change under randomization, is a white noise [92], and this is a property of the sequence of decimal digits of the sequence of digits of $\pi$. Randomization increases the Boltzmann entropy (Section 4.2 and equation (4–30) of [29]) or the algorithmic type [Sinai1959, 118, 42] of a random sequence; a sequence maximizes its entropy or equivalently, or has a maximum entropy or information content [Szilard1929, Shannon1948]. The conclusions could be falsified by assuming that the singularity in the infinite series functions used to calculate the digits [46] exist.
@@
For more details on the decimals of $\pi$ and its fractal dimension refer to the original work of Sevcik [Sevcik2017b], there. limited by the capacity of the available computers were calculated $10^{9}$ of $\pi$ with the algorithm of Bellard [23], already by that time the maximum number of decimal digits of $\pi$ known was $22.7\times 10^{1}$ 2 digits calculated with the same algorithm [23]. At the time of writing this book is known $2,699,999,990,000\approx 2,699\ldots\times 10^{12}$ decimals of $\pi$, [Pi2021, Sevcik2017b], store that number of decimals $\approx 2.7$ teraB of disk. But using $10^{9}$ of decimals of $\pi$ is enough to reach a value of $D_{S}\approx 1.888421\pm 10^{-6}$ while a sequence of a sequence of $10^{9}$ decimal digits of type $U_{\mathbb{Z}}(0.9)$ decimal has a $D_{S}\approx 1.88743881\pm 2\times 10^{-6}$.
@@
Since we do not know cial is the distribution of the decimals of $\pi$, the approximation of the inequality of Vysochanskij–Petunin [Vysochanskij1979, Vysochanskij1980, Vysochanskij1983] was used to compare the various values of $D_{S}$ obtained when the decimals of $\pi$ were increased by taking the first in the sequence $10,100,1000,\ldots,10^{7},10^{8},10^{9}$ digits of $\pi$ and na sequence of the same length of random digits distributed uniformly as $U_{\mathbb{Z}}(0.9)$ or $U_{\mathbb{R}}(0.1))$ versus the randomized sequence of the same number of the same digits of $\pi$. However, when the fractal dimension, evaluated as $D_{S}$, was used to compare the sequence of the decimals of $\pi$ against the same randomized sequence by, the sequence of the decimals of $\pi$ versus a sequence of the same length of decimals $U_{\mathbb{Z}}(0,9)$, no statistically significant differences were found ($0.460.11<P<$ 0.46). Pro comparing was compared with decimal sequences of $\pi$ with a length $\geqslant 100$ The sequence of real ($\in\mathbb{R}$) numbers which we call $U_{\mathbb{R}}(0.1)$ the $D_{S}$ was always statistically different from the $D_{S}$ of the sequence of the same number of decimals of $\pi$ [Sevcik2017b, Tables 1 and 2].
The price of approximating $D_{H}=\Phi$ by a simple way such as $D_{S}$ seems to require a very long sequence of data, we presume that the sequence of the decimals of $\pi$ is finally equal to that of a sequence of digits ($\in\mathbb{Z})$ which we call $U_{\mathbb{Z}}(0.9)$, both of the same length, $>10^{9}$. In general we think that $D_{S}=D_{H}=\Phi$ for sequences of $\aleph_{0}$ long digits of $\pi$, a way of saying that $D_{S}$ converges very slowly to its actual final value.
3.2.4 Multiple uses of $D_{S}$ the fractal dimension.
@@
The use of $D_{S}$ has increased considerably since its use for analyzing venoms was published (discussed below, [56, 59]), although perhaps the most important factor was that when Complexity International, where our work [Sevcik1998a] was originally published, ceased to be active and the work was deposited in www.arxive.org, [Sevcik2010]. With the use of www.arXiv.org, GoogleSholar starts tracking article usage (see https://bitly.ws/ZTbQ), but there are $D_{S}$ uses that go unreported in Scholar. We are aware of multiple attempts to calculate $D_{H}B$ and $D_{K}$ before 2010, some of which are here [72, 112, 70]
@@
Figure 3: Reverse phase high-performance liquid chromatography (HPLC) of Tityus discrepans. HPLC separation of the water-soluble venom T. discrepans was made with a column of C18 of reverse phase, the data were acquired at a frequency of 1 Hz, so that 60 min represent a $n=3600$ points. The cinematographic profile shown here corresponds to a batch milking of individual venoms, of 100 specimens of T. discrepans, put together as a lot poisons. For more details refer to the text of the book or the original publication [56].
@@
There are several situations where we consider complicated systems that we need to compare. An example appears when we need to compare the components that are separated in various chemical methods that are usually grouped under the name of chromatography. In these methods, a set of very diverse molecules are diluted in a liquid or gaseous medium that is forced to move through a solid medium, with which they interact and by that interaction, they separate forming “peaks” that are collected to study their properties.
@@
One of the chromatographic techniques with the greatest capacity to separate compounds is the so-called high-performance liquid chromatography, which separates components with different properties of electric charge, molecular weight or polarity [Snyder1979]. There is a huge number of examples of this analysis, here we will limit ourselves to consider the example of its use for the analysis of natural poisons produced by a genus of scorpions from South America, which we will arbitrarily limit to a genus called Tityus of which only in Venezuela more than 50 species [74] with great toxicity to humans are known.
@@
Figure 4: Two sources for calculating the fractal dimension of the elution pattern (${D}$ now ${D_{S}}$). The figure presents data on the venoms of the scorpions Tityus discrepans (panels on the left) and Rhopalurus laticauda (panels on the right). The chromatography protocol is the same as that used in Figure 3. The top panel presents the pattern of elution an extraction (milking) of each venom of an animal of each species as ordered. Data were acquired at a frequency of $1Hz$, which is equivalent to n=3600. The bottom panel presents the fractal dimension $D$ (now referred to as $D_{S}$) calculated for the complete stroke of the upper strokes and $D_{S}$ and was $1.17927\pm 8.01\times 10^{-5}$ [mean calculated with Eq. (10) and standard deviation as $\sqrt{{\text{var}\left[{(}\right]}{D_{S}}})$ calculated with Eq. (15) for $n=3600$ points of T. discrepans,and under the same conditions was $1.102637\pm 1.43\times 10^{-4}$, For R. laticauda. If you use a sliding window of $n$ points it is possible to calculate $Q$, assign that value to the center of the window and move it by 1 point, repeat the calculation of $Q$ and assign it to the next point in the abscissa, and so on until the available points of the waveform are exhausted. This is what is presented calculated with a sliding window of $n=500$ digitized points as explained for the upper panels, the value of $D_{S}$ obtained from the first $500$ points is assigned to the point $x=250$, the window is offset by 1 point, it is recalculated and the value is assigned to the point with $x=251$, and so on to the point $N-250$. For more details refer to Figure 3, the text of the book or the original publication [56].
@@
Here we will limit ourselves to consider the case of more impact, Tityus discrepans that coexists with the largest city in the country, Carcas of 4 million inhabitants with about 5000 annual cases registered [56, 57, 58, 59], but which is only one of the localities affected by scorpion ism by Tityus in Venezuela. From the venom of T. discrepans alone, some 206 fractions of toxic peptides [21] have been separated. The problem of the complexity of these poisons is highly relevant. To separate this large number of compounds, it is usually required to rechromatograph (repeat chromatography) under modified conditions the compounds obtained under the initial conditions.
@@
Figure 3 presents the initial result of a chromatography of a batch composed of the milking of venom of 100 specimens of T. discrepans. A large number of chromatographic peaks are observed, and different peaks often have different effects, even peaks that appear together in the chromatography profile usually have similar effects [58]. With elution profiles of such complexity, comparing poisons is complicated. And the complexity increases if one considers that the cinematographic profile varies between individuals of the same species, and varies seasonally as well.
@@
The Figure 4 presents the elution patterns of an individual milking of T. discrepans venom and an individual milking of venom of R. laticuauda, a species of scorpion common in areas of Venezuela below 500 meters above sea level, which poses no risk to humans. The figure is an example that scorpions can be compared from individual to individual with HPLC, so an instrument for comparing individual chromatography is very useful. The upper panels of Figure 4 show the chromatograms, T. discrepans, above and on the left, and R. laticauda, at the top right. Below each chromatogram is a graph of $D_{S}$ (formerly $D$) calculated with a sliding window of 500 points in length, as indicated in the text of the figure.
@@
Figure 5: Tortuosity graph constructed with a sliding window (VDT) $\boldsymbol{{Q}}$, where $\boldsymbol{{Q=D_{s}-1}}$. In cases such as the elution profiles of poisons, the waveform to be analyzed is obviously not infinite. In such cases fractal dimension can be questioned, and it is preferable to use the ‘tortuosity’ component of the fractal dimension that exceeds that of a line. $Q=D_{S}-1$. If a sliding sliding window of $n=500$ is used as explained for Figure 4, but here the set of $\{Q_{1,i}\}$ is plotted against the set $\{Q_{2,i}\}$, where $1$ expresses one kind of scorpion and $2$ the other; If both species were identical, the graph would be a line with slope $1$ and intercept 00. The figure shows that T. caripitensis and T. clathratos have similar venoms, but that the venom of Broteas libinally is very different from that of T. clathratus). For more details refer to Figure 4, the text of the book or the original publication [56].
@@
The Figure 5 is an example of using $D_{S}$ curves, such as those shown in the lower panels of the Figure 4, presented here. The figures were prepared by pairing point by point $\{D_{S_{i},1}\}_{i=1,\ldots,N}$ with the points of $\{D_{S_{i},2}\}_{i=1,\ldots,N}$, which was used co abscissa and which as ordered, was an arbitrary decision of the experimenter, it is easy to conceive that if $\{D_{S_{i},1}\}_{i=1,\ldots,N}=\{D_{S_{i},2}\}_{i=1,\ldots,N}$, the points on the graph would form a diagonal with slope 1. In the left panel of the figure, the sequence of $D_{S}$, T. clathratus was used (arbitrarily) as ordered and $D_{S}$, T. caripitensis as abscissa, although the correlation between the venoms is not perfect, the points are grouped around the line with slope 1 and intercept 0, it is good enough to reduce a coefficient of determination of 0.813. The graph on the left also has the $D_{S]}$, T. caripitensis as abscissa, but the ordinate corresponds to the venom of Brotheas libinalli $D_{S}$, B. libinalli as ordered; The distribution of the points in the graph is far from its diagonal. The venom of B. libinalli offers no danger to humans. The graph suggests that the venoms of T. clathratus and T. caripitensis are very similar, and that the poisons of B. libinalli and T. caripitensis are very different. The reader interested in more detail should refer to the urinals studies [56, 59].
@@
There is an indefinite number of systems that provide services to an indefinite number of service seekers. Indefinite, here, has the same implication of $\aleph_{0}$, as we have mentioned numerous times throughout this text. Some services (the bus that passes at a certain time through my stop, the emergency of a hospital, the supermarket or pharmacy where I buy, etc.) are very obviously service providers to those who come to them; In all of them there is an obvious service provider, and a queue or line, of those waiting to be served. Perhaps less obvious as a service are public traffic vans such as so-called highways. However, in all these cases the flow of those served follows a common function: a random Brownian march [34, 33, Samuel2012, 157, 152].
Four years ago I was consulted at the Madrid Hospital La Fuenfría, a hospital of the Madrid Health Service (SERMANS) dedicated to treating chronic cases. The reason for the consultation was to increase the efficiency of the Hospital, prone to prolonged periods of low occupancy. Full data was published [RodriguezHernandez2022] and refers interested readers to that publication for details omitted here. For the following analysis, daily Hospital occupancy data from May 1, 2014 to December 19, 2017 [RodriguezHernandez2022, corrections for asymmeter errors as indicated here].
Figure 6: Patients hospitalized every day (InP) at La Fuenfría Hospital from May 1, 2014 to December 19, 2017. The panels are: Panel A- Number of hospitalized patients (InP) Hospital La Fuenfría on each day of the Hospital study; Panel B- Daily variations (DInP) of patients staying at Hospital [$\Delta and$ in Eq. (19)]; Panel C- Panel A data expressed as fumtion of density of probability (fdp), the probability of observing a measurable or countable random event) for any number of patients (InP) who remain hospitalized each day during the study period ($\bullet$) [Wilks1962], a Gaussian line of type ($N[\bar{y},s[y]$) with mean $\bar{y}$ and variance ($s[y]^{2}$) both calculated for the 1329 data in Panel A ($\bar{y}=149.2$ and $s[y]=24.4$); Panel D- Daily differences between admissions (Figure 7A) and hospital discharges (Figure 6C) at the Hospital (DDiAd). Panel E- Random march $\boldsymbol{{\Xi_{i}}}$ built for DDiAd (Panel D) using Eq. (20), please note that the end resembles the InP sequence in Panel A. In all cases abscissa is the number of days between May 1, 2014. The largest negative or negative peaks are off-scale from the charts.
@@
Here the $y_{i}$ are daily data and the $\Delta y_{i}$ are the daily changes, of the various occupations of patients in the Hospital. The Figure 6A shows the sequence $\Delta y_{t}$ calculated with the Eq. (19), the curve covers a total of 1216 days of study, where viewed freely. During the study period, the number of hospitalized patients was $150;(149-151)$ (median, and a 95% CI, $n=$1329 days) in a range of ($102-187$) patients, meaning that Hospital occupancy is $0.77\;(0.53-0.77)$ (median, and, its 95% CI). There could be one 4 “cycles” where the curve goes from 60%? from occupancy to a total of 98% occupancy in a “cycles” of about 304? days ($\approx 10$ months) duration. If we accept those “cycles” as real, they do not follow any annual seasonal period that we could identify. something that is, at least, rare.
The Figure (6)C is a graph of the pdf of the daily variations of the data in the Figure (6)A, there it is observed that the data are not Gaussian, there are no negative data and there is a skewness towards positive high values.
Trying to understand the Figure 6A we build the Figures 6B and 6D. The first of these (Figure 6B) is the number of patients the Hospital receives daily, and the second (Figure 6D) is the number of patients the Hospital discharges (patients leaving the Hospital daily). From the simple observation of the data in the Figures 6B and 6D, there does not seem to be a pattern in them, they look like two “white noises” [92], random oscillations around a mean, which are commonly distributed in Gaussian form.
@@
It is surprising, however, that if we calculate the daily difference between income and expenses and graph them, we obtain the Figure 6E. This figure is constructed as the $\Xi_{i}$ of the Figure 6E, Built with the Eq. (20):
$\boldsymbol{{\Xi_{i}}}=\begin{cases}0&\implies i=1\\ DDiAd_{i}+\boldsymbol{{\Xi_{i--1}}}&\implies 2\leqslant i\leqslant 1328\;.\end{cases}$ (20)
Surprisingly, with the exception of a few peaks that go out of the curve, probably resulting from the subtractive cancellation of small differences, the Figure 6E is almost identical to the Figure 6A. The fundamental difference between the curves in the Figures 6B and D is that they reflect a delay in admitting a patient to the Hospital after another has been discharged, this shows the original work of Rodríguez-Hernández and Sevcik [RodriguezHernandez2022, Check Figure 4]. The Figures 6E and 6A are practically identical, and belie the idea that the apparent “cycles” are due to some seasonal factor external to La Fuenfría Hospital or the SERNAS hospital network of which the Hospital is a part. A.un more, the data favor the hypothesis that the oscillations of the occupation of the Hospital La Fuenfría, is due to factors that determine a delay in admitting a patient when another is discharged, not to some hidden factor external to the Hospital, or in the best of cases to the hospital network of SERMAS.
In the Figure 7 it becomes more evident that income and discharges have a positive median (Figure 7B and D) when a few negative data likely resulting from subtractive cancellation between small data are deleted (Figures 7A and 7C).
Figure 8: Two white boise sequences generated with Monte Carlo simulations [Panels A and C], as indicated by Box and Muller [31, 170]. The Panels are: Panel A, Gaussian White Noise and B: Gaussian Brownian Noise Rw, Rw generated using Eq. (23) forms the data in panel A;Panel C, Gaussian white noise and $\{g_{i}=\text{{{N}}}\left\{0,10^{-4}\right\}\}_{i=1,2,\ldots,10^{4}}$; D: Rw Gaussian, a Rw generated using the Eq. (23) with panel C data, the abscissa and ordinate are arbitrary. Please note that $\left[D_{S}\pm s(D_{S})\right]$ estimated with Eqs. (10) and (15) are: Panel A, $1.66122\pm 7.60191\cdot 10^{-4}$; Panel B, $1.32692\pm 7.6926\cdot 10^{-4}$; Panel C, $1.6692\pm 7.64176\cdot 10^{-4}$; Panel D, $1.31723\pm 7.72806\cdot 10^{-4}$. Rw is a random walk, of the original work. See the original reference and text of this book for more data and details [RodriguezHernandez2022]
The relationship between Figures 6A and 6E can be understood if one uses Monte Carlo simulation to generate data around a known pdf and Marcovian time series [149, Shi2012]. In principle, we can visualize two kinds of time series where data that have a pdf appear such as $y_{i}=g(x_{i}\;\mid\;\{\theta\})$, where $\{\theta\}$ is a set of parameters on which $y_{i}$ depends. Then we can conceive of a series such as:
@@
In the first case any point $(y_{j}\perp\!\!\!\perp y_{i})_{\forall(i,j)}$, a waveform we call white noise. In the second case $(y_{i}\not{\perp\!\!\!\perp}y_{j})_{\forall(i,j)}$, but we cannot say how, this is what is called a Markovian process or a Markov series, the sequence that constitutes the waveform is usually called Brownian noise, since that kind of movements was discovered in the nineteenth century as characteristic of pollen grains in liquid medium by Brown [34, 35]. In white and Brown noises of the most common type, however, the pdf is Gaussian such that $g(x_{i}\;\mid\;\{\theta\})=\text{{{N}}}\left\{\mu,\sigma^{2}\right\}$, with mean $\mu$ and variance $\sigma^{2}$. Brownian noise follows a function as:

### v3

@@ Abstract
is obtained under randomization. I.e.: $D_{S}$ does not change under randomization, is a white noise [94], and this is a property of the sequence of decimal digits of the sequence of digits of $\pi$. Randomization increases the Boltzmann entropy (Section 4.2 and equation (4–30) of [29]) or the algorithmic type [Sinai1959, 120, 43] of a random sequence; a sequence maximizes its entropy or equivalently, or has a maximum entropy or information content [Szilard1929, Shannon1948]. The conclusions could be falsified by assuming that the singularity in the infinite series functions used to calculate the digits [47] exist.
@@
For more details on the decimals of $\pi$ and its fractal dimension refer to the original work of Sevcik [Sevcik2016, Sevcik2017b], there. limited by the capacity of the available computers were calculated $10^{9}$ of $\pi$ decimal digits with the algorithm of Bellard [23], already by that time the maximum number of decimal digits of $\pi$ known was $2.7\times 10^{12}$ digits calculated with the same algorithm [23]. At the time of writing this article $2,699,999,990,000\approx 2,699\ldots\times 10^{12}$ decimals of $\pi$ are known, [12], store that number of decimals requires $\approx 2.7$ teraB of disk. But using $10^{9}$ of decimals of $\pi$ is enough to reach a value of $D_{S}\approx 1.888421\pm 10^{-6}$ while a sequence of of $10^{9}$ decimal digits of type $U_{\mathbb{Z}}(0.9)$ decimal has a $D_{S}\approx 1.88743881\pm 2\times 10^{-6}$.
@@
Since we do not know the distribution of the decimals of $\pi$, the approximation of the inequality of Vysochanskij–Petunin [Vysochanskij1979, Vysochanskij1980, Vysochanskij1983] was used to compare the diverse values of $D_{S}$ obtained when the decimals of $\pi$ studied were increased by taking the first in the sequence of $10,100,1000,\ldots,10^{7},10^{8},10^{9}$ $\pi$ digits and a sequence of the same length of decimal distributed uniformly as $U_{\mathbb{Z}}(0.9)$ or $U_{\mathbb{R}}(0.1))$ versus the randomized sequence of the same number of $\pi$ digits. However, when $D_{S}$, was used to compare the sequence o $\pi$ decimals with the same randomized sequence of $\pi$ decimals and a sequence of the same length of $U_{\mathbb{Z}}(0,9)$ decimals, no statistically significant differences were found ($0.460.11<P<$ 0.46).
3.3 Other examples of the uses of $D_{S}$.
@@
$D_{S}$ used to analyze venoms was published (discussed below, [57, 60]), although perhaps the most important factor was that when Complexity International, where our work [Sevcik1998a] was originally published, ceased to be active and the work was deposited in www.arXive.org, [Sevcik2010]. With the use of www.arXiv.org, GoogleSholar starts tracking article usage (see https://scholar.google.com/citations?hl=en&user=0VhfpTIAAAAJ), but there are $D_{S}$ uses that are unreported in GoogleScholar. We are aware of multiple attempts to calculate $D_{HB}$ and $D_{K}$ before 2010, some of which are discussed here [74, 114, 72]
@@
Figure 3: Reverse phase high-performance liquid chromatography (HPLC) of Tityus discrepans scorpion venom. HPLC separation of the water-soluble venom T. discrepans was made with a column of C18 of reverse phase, the data were acquired at a frequency of 1 Hz, so that 60 min represent a $n=3600$ points. The cknematographic profile shown here corresponds to a batch milking of individual venoms, of 100 specimens of T. discrepans, put together as a lot poisons. For more details refer to the text of the book or the original publication [57].
@@
There are several situations where we consider complicated [^3] The word “complicated” is used here to re preserve “complex” for numbers of type $(a+b\sqrt{-1})\in\mathbb{C}$. systems that we need to compare. An example appears when we need to compare the components that are separated in various chemical methods that are usually grouped under the name of chromatography. In these methods, a set of very diverse molecules are diluted in a liquid or gaseous medium that is forced to move through a solid medium, with which they interact and by that interaction, they separate forming “peaks” that are collected to study their properties.
@@
One of the chromatographic techniques with the greatest capacity to separate compounds is the so-called high-performance liquid chromatography (HPLC), which separates components with different properties of electric charge, molecular weight or polarity [Snyder1979]. There is a huge number of examples of this analysis, here we will limit ourselves to consider the example of its use for the analysis of natural poisons (called ‘toxins’), produced by a genus of scorpions from South America, which we will arbitrarily limit to a genus called Tityus of which only in Venezuela more than 50 species [76] with great toxicity to humans are known.
@@
Figure 4: Two sources for calculating the fractal dimension (${D_{S}}$) of the elution pattern . The figure presents data on the venoms of the scorpions Tityus discrepans (panels on the left) and Rhopalurus laticauda (panels on the right). The chromatography protocol is the same as that used in Figure 3. The top panel presents the pattern of elution an extraction (milking) of each venom of an animal of each species as ordered. Data were acquired at a frequency of $1Hz$, which is equivalent to n=3600. The bottom panel presents the fractal dimension $D_{S}$ calculated for the upper traces and $D_{S}$ was $1.17927\pm 8.01\times 10^{-5}$ [mean calculated with Eq. (10) and standard deviation as $\sqrt{{\text{var}\left[{(}\right]}{D_{S}}})$ calculated with Eq. (15) for $n=3600$ points of T. discrepans,and under the same conditions was $1.102637\pm 1.43\times 10^{-4}$, For R. laticauda. If you use a sliding window of $n$ points it is possible to calculate $Q=D_{s}-1$, assign that value to the center of the window and move it by 1 point, repeat the calculation of $Q$ and assign it to the next point in the abscissa, and so on until the available points of the waveform are exhausted. This is what is presented calculated with a sliding window of $n=500$ digitized points as explained for the upper panels, the value of $D_{S}$ obtained from the first $500$ points is assigned to the point $x=250$, the window is offset by 1 point, it is recalculated and the value is assigned to the point with $x=251$, and so on to the point $N-250$. For more details refer to Figure 3, the text of the article or the original publication [57].
@@
Here we will limit ourselves to consider the case of more impact, Tityus discrepans that coexists with the largest city in the country, Carcas of 4 million inhabitants with about 5000 annual scorpion sting cases registered [57, 58, 59, 60], but which is only one of the localities affected by scorpion ism by Tityus in Venezuela. From the venom of T. discrepans alone, some 206 fractions of toxic peptides [21] have been separated. The problem of the complexity of these poisons is highly relevant. To separate this large number of compounds, it is usually required to rechromatograph (repeat chromatography) under modified conditions the compounds obtained under the initial conditions.
@@
Figure 3 presents the initial result of a chromatography of a batch composed of the milking of venom of 100 specimens of T. discrepans. A large number of chromatographic peaks are observed, and different peaks often have different effects, peaks that appear together in the chromatography profile usually have similar effects [59]. With elution profiles of such complexity, comparing poisons is complicated. And the complexity increases if one considers that the chromatographic profile varies between individuals of the same species, and varies seasonally as well.
@@
Figure 4 presents the elution patterns of an individual (meang 1 milking from 1 scorpion)) milking of T. discrepans venom and an individual milking of venom of R. laticuauda, a species of scorpion common in areas of Venezuela below 500 meters above sea level, which poses no risk to humans. The figure is an example that scorpions can be compared from individual to individual with HPLC, so an instrument for comparing individual chromatography is very useful. The upper panels of Figure 4 show the chromatograms, T. discrepans, above and on the left, and R. laticauda, at the top right. Below each chromatogram is a graph of $D_{S}$ calculated with a sliding window of 500 points in length, as indicated in the text of the figure.
@@
Figure 5: Tortuosity graph constructed with a sliding window $\boldsymbol{{Q=D_{s}-1}}$. In cases such as the elution profiles of poisons, the waveform to be analyzed is obviously not infinite. In such cases fractal dimension can be questioned, and it is preferable to use the ‘tortuosity’ component of the fractal dimension that exceeds that of a line. If a sliding sliding window of $n=500$ is used as explained for Figure 4, but here the set of $\{Q_{1,i}\}$ is plotted against the set $\{Q_{2,i}\}$, where $1$ expresses one kind of scorpion and $2$ the other; If both species were identical, the graph would be a line with slope $1$ and intercept 00. The figure shows that T. caripitensis and T. clathratos have similar venoms, but that the venom of Broteas libinally is very different from that of T. clathratus). For more details refer to Figure 4, the text of the book or the original publication [57].
@@
Figure 5 is an example of using $Q$ curves, such as those shown in the lower panels of the Figure 4. The figures were prepared by pairing point by point $\{Q_{i,1}\}_{i=1,\ldots,N}$ with the points of $\{Q_{i,2}\}_{i=1,\ldots,N}$ as ordinate. Which of the two sets was the abscissa and which ordinate, were arbitrary decisions of the experimenter. It is easy to conceive that if $\{Q_{i,1}\}_{i=1,\ldots,N}$ and $\{Q_{i,2}\}_{i=1,\ldots,N}$ were equal, the points on the graph would form a diagonal with slope 1. In the left panel of the figure, the sequence for T. clathratus was used as ordered, and T. caripitensis as abscissa. Although the correlation between the venoms is not perfect, the points are grouped around the line with slope 1 and intercept 0, is good enough to produce a coefficient of determination of 0.813. The graph on the left also has T. caripitensis as abscissa, but the ordinate corresponds to the venom of Brotheas libinalli $D_{S}$. The distribution of the points in the graph is far from its diagonal. The venom of B. libinalli offers no danger to humans. The graph suggests that the venoms of T. clathratus and T. caripitensis are very similar, and that the venoms of B. libinalli and T. caripitensis are very different. The reader interested in more detail should refer to the urinals studies [57, 60].
@@
There is an indefinite number of systems that provide services to an indefinite number of service seekers. Indefinite, here, has the same implication of $\aleph_{0}$, an indefinitely large number fulfilling $\aleph_{0}\in\mathbb{Z}$. Some services (the bus that passes at a certain time through my stop, the emergency of a hospital, the supermarket or pharmacy where I buy, etc.) are very obviously service provided to those demanding them. In all of them there is an obvious service provider, and a queue or line, of those waiting to be served. Perhaps less obvious as a service are public traffic means such as the, so-called highways. However, in all these cases the flow of those served follows a common function: a random Brownian march [35, 34, Samuel2012, 159, 154].
Four years ago I was consulted at the La FuenfríaMadrid Hospital, a hospital of the Madrid Health Service (SERMANS) dedicated to treating chronic cases. The reason for the consultation was to increase the efficiency of the Hospital, prone to prolonged periods of low occupancy. Full data was published [RodriguezHernandez2022] and we refer interested readers to that publication for details omitted here. For the analysis, daily Hospital occupancy data from May 1, 2014 to December 19, 2017 [RodriguezHernandez2022, corrections for asymmeter errors as indicated here].
Figure 6: Patients hospitalized every day (InP) at La Fuenfría Hospital from May 1, 2014 to December 19, 2017. The panels are: Panel A- Number of hospitalized patients (InP) at Hospital La Fuenfría on each day of the Hospital study; Panel B- Daily variations (DInP) of patients staying at Hospital [$\Delta and$ in Eq. (19)]; Panel C- Panel A data expressed as fumtion of density of probability (fdp), the probability of observing a measurable or countable random event) for any number of patients (InP) who remain hospitalized each day during the study period ($\bullet$) [Wilks1962], a Gaussian line of type ($N[\bar{y},s[y]$) with mean $\bar{y}$ and variance ($s[y]^{2}$) both calculated for the 1329 data in Panel A ($\bar{y}=149.2$ and $s[y]=24.4$); Panel D- Daily differences between admissions (Figure 7A) and hospital discharges (Figure 6C) at the Hospital (DDiAd). Panel E- Random march $\boldsymbol{{\Xi_{i}}}$ built for DDiAd (Panel D) using Eq. (20), please note that the end resembles the InP sequence in Panel A. In all cases abscissa is the number of days between May 1, 2014. The largest negative or negative peaks are off-scale from the charts.
@@
Here the $y_{i}$ are daily data and the $\Delta y_{i}$ are the daily changes, of the various occupations of patients in the Hospital. The Figure 6A shows the sequence $\Delta y_{t}$ calculated with the Eq. (19), the curve covers a total of 1216 days of study. During the study period, the number of hospitalized patients was $150\;(149-151)$ (median, and a 95% CI, $n=$1329 days) in a range of ($102-187$) patients, meaning that Hospital occupancy is $0.77\;(0.53-0.77)$ (median, and, its 95% CI). There could be one 4 “cycles” where the curve goes from 60% from occupancy to a total of 98% occupancy in a “cycles” of about 304 days ($\approx 10$ months) duration?. If we accept those “cycles” as real, they do not follow any annual seasonal period that we could identify. something that is, at least, rare.
Figure (6)C is a graph of the pdf of daily variations of the data in Figure (6)A, there it is observed that the data are not Gaussian, there are no negative data and there is a skewness towards positive high values.
Trying to understand Figure 6A we built Figures 6B and 6D. The first of these (Figure 6B) is the number of patients the Hospital receives daily, and the second (Figure 6D) is the number of patients the Hospital discharges (patients leaving the Hospital) daily. From the simple observation of the data inFigures 6B and 6D, there does not seem to be a pattern in them, they look like two “white noises” [94], random oscillations around a mean, which are commonly distributed in Gaussian form.
@@
It is surprising, however, that if we calculate the daily difference between input and output and graph them, we obtain the Figure 6E. This figure is constructed as the $\Xi_{i}$ of the Figure 6E, Built with Eq. (20):
$\boldsymbol{{\Xi_{i}}}=\begin{cases}0&\implies i=1\\ DD_{i}Ad_{i}+\boldsymbol{{\Xi_{i--1}}}&\implies 2\leqslant i\leqslant 1328\;.\end{cases}$ (20)
Surprisingly, with the exception of a few peaks that go out of the curve, probably resulting from the subtractive cancellation of small differences, the Figure 6E is almost identical to the Figure 6A. The fundamental difference between the curves in the Figures 6B and D is that they reflect a delay in admitting a patient to the Hospital after another has been discharged, this shows the original work of Rodríguez-Hernández and Sevcik [RodriguezHernandez2022, Check Figure 4]. The Figures 6E and 6A are practically identical, and belie the idea that the apparent “cycles” are due to some seasonal factor external to La Fuenfría Hospital or the SERNAS hospital network of which the Hospital is a part. And more, the data favor the hypothesis that the oscillations of the occupation of the Hospital La Fuenfría, is due to factors that determine a delay in admitting a patient when another is discharged, not to some hidden factor external to the Hospital, or in the best of cases to the hospital network of SERMAS.
In Figure 7 it becomes more evident that income and discharges have a positive median (Figure 7B and D) when a few negative data likely resulting from subtractive cancellation between small data are deleted (Figures 7A and 7C).
Figure 8: Two white boise sequences generated with Monte Carlo simulations [Panels A and C], as indicated by Box and Muller [31, 172]. The Panels are: Panel A, Gaussian White Noise and B: Gaussian Brownian Noise Rw, generated using Eq. (23) forms the data in panel A;Panel C, Gaussian white noise and $\{g_{i}=\text{{{N}}}\left\{0,10^{-4}\right\}\}_{i=1,2,\ldots,10^{4}}$; D: Rw Gaussian, generated using the Eq. (23) with panel C data, the abscissa and ordinate are arbitrary. Please note that $\left[D_{S}\pm s(D_{S})\right]$ estimated with Eqs. (10) and (15) are: Panel A, $1.66122\pm 7.60191\cdot 10^{-4}$; Panel B, $1.32692\pm 7.6926\cdot 10^{-4}$; Panel C, $1.6692\pm 7.64176\cdot 10^{-4}$; Panel D, $1.31723\pm 7.72806\cdot 10^{-4}$. Rw is a random walk, of the original work. See the original reference and text of this book for more data and details [RodriguezHernandez2022]
The relationship between Figures 6A and 6E can be understood if one uses Monte Carlo simulation to generate data around a known pdf and Marcovian time series [151, Shi2012]. In principle, we can visualize two kinds of time series where data that have a pdf appear such as $y_{i}=g(x_{i}\;\mid\;\{\theta\})$, where $\{\theta\}$ is a set of parameters on which $y_{i}$ depends. Then we can conceive of a series such as:
@@
In the first case any point $(y_{j}\perp\!\!\!\perp y_{i})_{\forall(i\neq j)}$444$\forall$ means ‘for all’, is a waveform we call white noise. In the second case $(y_{i}\not{\perp\!\!\!\perp}y_{j})_{\forall(i\neq j)}$, but we cannot say how, this is what is called a Markovian process or a Markov series, the sequence that constitutes the waveform is usually called Brownian noise, since that kind of movement was discovered in the nineteenth century as characteristic of pollen grains in liquid medium by Brown [35, 36]. In white and Brown noises of the most common type, however, the pdf is Gaussian such that $g(x_{i}\;\mid\;\{\theta\})=\text{{{N}}}\left\{\mu,\sigma^{2}\right\}$, with mean $\mu$ and variance $\sigma^{2}$. Brownian noise follows a function as: